# Feshbach-Villars Formalism for a Spin-1/2 Particle in Curved Spacetime

### v1

@@
Abdelmalek Boumali abdelmalek.boumali@univ-tebessa.dz or boumali.abdelmalek@gmail.com Laboratory of theoretical and applied Physics,
Echahid Cheikh Larbi Tebessi University, Algeria
(April 1, 2025)
This study explores the Feshbach-Villars (FV) formalism for spin-1/2 particles in curved spacetime. We derive the Hamiltonian form of the Dirac equation in this context and extend the FV transformation accordingly. The generalized Klein-Gordon equation, obtained by squaring the Dirac operator, is reformulated using the FV approach. We present the resulting Hamiltonian in both matrix and Pauli matrix forms, considering gravitational and electromagnetic interactions. The formalism is examined in both (1+2) and (1+3) dimensional spacetimes, with special attention given to the spin-field interaction term. This study provides a framework for studying relativistic quantum mechanics in curved spacetime, offering insights into the interplay between quantum effects, gravity, and electromagnetism.
@@ I Introduction
Relativistic quantum mechanics represents a significant academic achievement that focuses on integrating the concepts of relativity and quantum mechanics. This integration expanded the framework of quantum theory and introduced the concept of negative energy levels, which emerged directly from relativistic formulations. Negative energy states are illustrated by the concept of the Dirac sea, which is an infinite expanse of particles with negative energy. In this model, the absence of a particle in the Dirac sea indicates the presence of a physical particle, specifically, a positron. This interpretation arises from the requirement that the energy of free particles must be positive, as dictated by the physical laws. By treating particles and antiparticles differently, this approach creates intrinsic asymmetry. While this asymmetry is essential for upholding the probabilistic principles of quantum physics, as outlined in the continuity equation, it also highlights the elegance of Dirac’s sea model. This concept is particularly applicable to fermions, which have a half-integer spin and obey the Pauli exclusion principle. However, this does not apply to bosons, which possess an integer spin. Quantum field theory (QFT) provides a more comprehensive framework by treating antiparticles as fundamental components of field interactions rather than relying on the Dirac sea [1, 2, 3].
@@
In 1958, Feshbach and Villars [1] published their findings on linearizing the Klein-Gordon equation. Their work led to the development of the Feshbach-Villars (FV-0) equations, which allowed the Klein-Gordon equation to be expressed in the Schrödinger form, also known as the Hamiltonian form. This formalization addresses various contradictions in the Klein-Gordon equation, including the existence of negative energies. This facilitated a clearer interpretation of probability theory, particularly when considering particles with corresponding antiparticles.
@@
Feshbach and Villars’ formalism resolved the key issues that had previously caused the abandonment of the Klein-Gordon theory, enabling a reevaluation of its principles. Significant research has been conducted using the Feshbach-Villars equations. Furthermore, efforts have been made to obtain numerical solutions and approximations for these equations, providing physicists with the means to tackle ongoing challenges and develop insights related to relativistic quantum mechanics (RQM) [4, 5, 6, 7, 8, 9, 10].
@@
To expand the theory within the framework of general relativity, Silenko [11, 12] was the first to introduce a new form of conformal invariance for massless scalar particles in Riemannian spacetimes. He studied scalar particles in arbitrary spacetimes under general relativity (GR). To derive a Hamiltonian form of the covariant Klein-Gordon (KG) equation for both massive and massless particles, he utilized the generalized Feshbach-Villars transformation, as proposed in a previous study. Silenko then derived the general quantum-mechanical equations of motion and established their classical limits. Recently, this method has been applied in various contexts, and the results have been promising [13, 14, 15, 16].
@@
Building on this idea, FV-like equations have been proposed for spin-½ particles (KG–½). These equations have the potential to achieve complete particle-antiparticle symmetry and eliminate the need for a Dirac sea, representing a promising development. By altering the manner in which wave functions depict physical states, this extension alters certain scientific interpretations, particularly those involving transition probabilities. The differences become significant when examining phenomena where the influence of the Dirac sea is evident. The formalism presented by Feshbach and Villars paved the way for a new relativistic quantum mechanics equation for spin-1/2 particles. The worked developed by Robson and Staudte [2, 3] followed the idea of the Feshbach-Villars linearization procedure and applied it to the second-order Dirac equation. The result was an eight-component relativistic equation in Hamiltonian form, which preserves some of the inner product properties and the probabilistic interpretation obtained from the FV-0 equation. This eight-component equation or Feshbach-Villars equation for spin- 1 / 2 particles (FV-1/2) is relativistic and different from the Dirac equation.
@@
This study aims to extend the Feshbach-Villars (FV) formalism for spin-1/2 particles to curved spacetime. Specifically, we analyze fermion particles in arbitrary spacetimes within the framework of general relativity (GR). To derive the Hamiltonian form of the covariant Dirac equation in curved spacetime, we employ the generalized Feshbach-Villars transformation introduced in Ref. [12]. It is worth noting that, to the best of our knowledge, there is a notable gap in the literature regarding this type of study.
II Hamiltonian form of Dirac equation in curved spacetime
@@
. (4)
Multiply (2) through by$(\gamma^{0})^{-1}$ (note: in most conventions, $(\gamma^{0})^{-1}=\gamma^{0}$), give
@@
  • $-i\gamma^{0}\gamma^{i}\mathcal{D}_{i}$: Represents the kinetic energy and coupling of the spinor to the spatial curvature via the spin connection $\omega_{i}^{ab}$.
@@
  • $-\frac{1}{4}\gamma^{0}\omega_{0}^{ab}\gamma_{ab}$: Represents the coupling to the spin connection in the time direction.
@@
where $\nabla_{\mu}$ is the covariant derivative compatible with the curved spacetime metric. Thus, the probability density $\rho$ is the 00-component of the current:
@@
III Feshbash-Villars of a particle of spin-1/2 in curved spacetime: The formalism
The modified Klein-Gordon equation obtained by squaring the operator in the Dirac equation, first found by Erwin Schrödinger is given by
@@ III Feshbash-Villars of a particle of spin-1/2 in curved spacetime: The formalism
Following the procedure used by Silenko (12, 2025) the general form of the covariant KG equation reads(13) (see Annex A)
@@
[…]
@@
Using the FV transformation, the Klein-Gordon equation is rewritten as a first-order equation for $\psi$.
Substituting the definitions of $\phi$ and $\chi$, we derive a coupled system of equations:
@@
$-e\vec{\Sigma}\cdot\vec{B}=-e\begin{pmatrix}\vec{\sigma}&0\\ 0&\vec{\sigma}\end{pmatrix}\cdot\vec{B}.$ (43)

### v2

@@
Abdelmalek Boumali abdelmalek.boumali@univ-tebessa.dz or boumali.abdelmalek@gmail.com Echahid Cheikh Larbi Tebessi University, Tebessa, Algeria
(February 24, 2026)
This study extends the Feshbach–Villars (FV) formalism to spin–$1/2$ particles propagating in curved spacetime. Starting from the covariant Dirac equation, we derive its Hamiltonian form in a tetrad (vierbein) formulation and implement a diagonal FV transformation of the associated second-order (squared) equation. The resulting Schrödinger-type evolution equation is written in both matrix and Pauli-matrix representations and explicitly displays gravitational and electromagnetic couplings, including spin–field interaction terms. We analyze the formalism in $(1+2)$- and $(1+3)$-dimensional settings and apply it to cosmic-string backgrounds (static and spinning, with and without oscillator couplings), obtaining the corresponding spectra and wave functions.
While the FV reformulation is algebraically equivalent to the covariant Dirac dynamics (up to the standard restriction that removes the spurious solutions introduced by squaring), it offers a complementary viewpoint that is useful in practice: (i) the particle/antiparticle sectors are represented by two coupled components with a simple (pseudo-)unitary inner product tied to the conserved charge; (ii) the effect of stationarity versus staticity is encoded directly in the FV operator $\mathrm{Y}$, which is controlled by the ADM shift vector and therefore diagnoses how frame–dragging backgrounds modify particle–antiparticle mixing; and (iii) the FV block structure provides a compact route to spectral problems in topologically nontrivial geometries.
@@ I Introduction
Relativistic quantum mechanics represents a significant academic achievement that focuses on integrating the concepts of relativity and quantum mechanics. This integration expanded the framework of quantum theory and introduced the concept of negative energy levels, which emerged directly from relativistic formulations. Negative energy states are illustrated by the concept of the Dirac sea, which is an infinite expanse of particles with negative energy. In this model, the absence of a particle in the Dirac sea indicates the presence of a physical particle, specifically, a positron. This interpretation arises from the requirement that the energy of free particles must be positive, as dictated by the physical laws. By treating particles and antiparticles differently, this approach creates intrinsic asymmetry. Although this asymmetry is essential to maintain the probabilistic principles of quantum physics, as described in the continuity equation, it also highlights the elegance of Dirac’s sea model. This concept is particularly applicable to fermions that have a half-integer spin and obey the Pauli exclusion principle. However, this does not apply to bosons that possess an integer spin. Quantum field theory (QFT) provides a more comprehensive framework by treating antiparticles as fundamental components of field interactions rather than relying on the Dirac sea Feshbach and Villars (1958); Robson and Staudte (1996); Staudte (1996) .
@@
In 1958, Feshbach and Villars Feshbach and Villars (1958) published their findings on linearizing the Klein-Gordon equation. Their work led to the development of the Feshbach-Villars (FV-0) equations, which allowed the Klein-Gordon equation to be expressed in the Schrödinger form, also known as the Hamiltonian form. This formalization addresses various contradictions in the Klein-Gordon equation, including the existence of negative energies. This facilitated a clearer interpretation of probability theory, particularly when considering particles with corresponding antiparticles.
@@
Feshbach and Villars’ formalism resolved the key issues that had previously caused the abandonment of the Klein-Gordon theory, enabling a reevaluation of its principles. Significant research has been conducted using the Feshbach-Villars equations. Furthermore, efforts have been made to obtain numerical solutions and approximations for these equations, providing physicists with the means to tackle ongoing challenges and develop insights related to relativistic quantum mechanics (RQM) Merad et al. (2000); Bounames and Chetouani (2001); Haouat and Chetouani (2005); Guettoul and Chetouani (2006); Brown et al. (2015); Motamedi et al. (2019); Wingard et al. (2024).
@@
To extend the FV programme to general relativistic settings, Silenko Silenko (2008, 2013) identified a conformal invariance of the massless scalar theory in Riemannian spacetimes, i.e. an invariance of the properly defined wave equation under local Weyl rescalings $g_{\mu\nu}\to\Omega^{2}(x)g_{\mu\nu}$ (with a corresponding field rescaling). This property is useful because it constrains the curvature-coupling term and allows one to construct a Hamiltonian form that behaves consistently under conformal transformations. He studied scalar particles in arbitrary spacetimes under general relativity (GR). To derive a Hamiltonian form of the covariant Klein-Gordon (KG) equation for both massive and massless particles, he utilized the generalized Feshbach-Villars transformation, as proposed in a previous study. Silenko then derived the general quantum-mechanical equations of motion and established their classical limits. Recently, this method has been applied in various contexts, and the results have been promising Bouzenada et al. (2023a); Garah and Boumali (2025b, a); Bouzenada et al. (2023b); Bouzenada and Boumali (2023); Bouzenada et al. (2024).
@@
Building on this idea, FV-like equations have been proposed for spin-½ particles (KG–½). These equations have the potential to achieve complete particle-antiparticle symmetry and eliminate the need for a Dirac sea, representing a promising development. By recasting the second-order Dirac dynamics into an enlarged FV state space, the wave function is represented by coupled components associated with particle and antiparticle sectors. This change modifies the underlying inner product (from a probability-density norm to a charge-density norm), and therefore affects how one interprets amplitudes and transition probabilities in the single-particle theory. The differences become significant when examining phenomena where the influence of the Dirac sea is evident. The formalism presented by Feshbach and Villars paved the way for a new relativistic quantum mechanics equation for spin-1/2 particles. The worked developed by Robson and Staudte Robson and Staudte (1996); Staudte (1996) followed the idea of the Feshbach-Villars linearization procedure and applied it to the second-order Dirac equation. The result was an eight-component relativistic equation in Hamiltonian form, which preserves some of the inner product properties and the probabilistic interpretation obtained from the FV-0 equation. This eight-component equation or Feshbach-Villars equation for spin- 1 / 2 particles (FV-1/2) is relativistic and different from the Dirac equation.
@@
This study aims to extend the Feshbach-Villars (FV) formalism for spin-1/2 particles to curved spacetime. Specifically, we analyze fermion particles in arbitrary spacetimes within the framework of general relativity (GR). To derive the Hamiltonian form of the covariant Dirac equation in curved spacetime, we employ the generalized Feshbach-Villars transformation introduced in Ref. Silenko (2013). It is worth noting that, to the best of our knowledge, there is a notable gap in the literature regarding this type of study.
It is important to stress the scope of the present work in light of the well-known fact that the FV construction starts from (and therefore remains anchored to) the covariant Dirac theory. The motivation is not to propose a new relativistic wave equation, but to provide a Hamiltonian formulation in an enlarged FV state space in which (a) the separation into positive-/negative-frequency sectors becomes explicit, (b) the conserved inner product is naturally expressed in terms of a simple metric operator, and (c) the role of geometric data (staticity, shift vector, and topological identifications) in the mixing between sectors can be tracked directly at the operator level. These points are particularly transparent in stationary defect geometries with nontrivial $g_{0\varphi}$ or $g_{z\varphi}$ components, where the FV operator $\mathrm{Y}$ does not vanish and the azimuthal quantum number is shifted by energy- and momentum-dependent terms (see Sec. V).
The paper is organized as follows. Section II reviews the FV linearization for the second-order Dirac equation in flat spacetime and establishes the notation used later. Section III derives the Hamiltonian form of the covariant Dirac equation in curved spacetime, emphasizing the tetrad fields and the spin connection. Section IV develops the FV formulation for spin–$1/2$ particles in general curved backgrounds and discusses the transformation properties of the resulting Hamiltonian. Section V applies the formalism to cosmic-string geometries (including oscillator couplings) and extracts the corresponding spectra. We conclude with a discussion of the scope and limitations of the approach and possible extensions.
II Solution of Relativistic oscillators of Feshbach-Villars of Spin-1/2
II.1 Theoretical framework
The Dirac equation for a charged particle in an electromagnetic field is Greiner (2001):
$\left(i\gamma^{\mu}D_{\mu}-m\right)\psi=0$ (1)
where $D_{\mu}=\partial_{\mu}+\frac{iq}{\hbar c}A_{\mu}$, and $A_{\mu}$ is the electromagnetic four-potential. Multiply both sides by $\left(i\gamma^{\nu}D_{\nu}+mc\right)$:
$\left(i\gamma^{\nu}D_{\nu}+mc\right)\left(i\gamma^{\mu}D_{\mu}-m\right)\psi=0$ (2)
Expanding this gives:
$\left[(i\gamma^{\nu}D_{\nu})(i\gamma^{\mu}D_{\mu})-m^{2}\right]\psi=0$ (3)
or
$\gamma^{\nu}\gamma^{\mu}D_{\nu}D_{\mu}\psi-m^{2}\psi=0$ (4)
The gamma matrices satisfy the anticommutation relation:
$\{\gamma^{\mu},\gamma^{\nu}\}=2g^{\mu\nu}$ (5)
So,
$\gamma^{\nu}\gamma^{\mu}=g^{\nu\mu}+\sigma^{\nu\mu}$ (6)
where
$\sigma^{\nu\mu}=\frac{1}{2}[\gamma^{\nu},\gamma^{\mu}]$ (7)
Therefore,
$\gamma^{\nu}\gamma^{\mu}D_{\nu}D_{\mu}=-g^{\nu\mu}D_{\nu}D_{\mu}+-\sigma^{\nu\mu}D_{\nu}D_{\mu}$ (8)
The first term gives the d’Alembertian (wave operator), the second term gives the coupling to the electromagnetic field tensor $F_{\mu\nu}$ via the commutator:
$D_{\nu},D_{\mu}]=iqF_{\nu\mu},\,F_{\nu\mu}=\partial_{\nu}A_{\mu}-\partial_{\mu}A_{\nu}$ (9)
So,
$\sigma^{\nu\mu}D_{\nu}D_{\mu}=q\sigma^{\nu\mu}F_{\nu\mu}$ (10)
Putting it all together:
$\left[D^{\mu}D_{\mu}-m^{2}+q\sigma^{\nu\mu}F_{\nu\mu}\right]\psi=0$ (11)
The operator $D^{\mu}D_{\mu}$ is the covariant d’Alembertian, and $\sigma^{\nu\mu}F_{\nu\mu}$ gives the coupling to the electromagnetic field, which can be written in terms of electric and magnetic fields using the Pauli matrices and their combinations.
The Dirac equation for a charged spin-1/2 particle in an electromagnetic field is "squared" to obtain a second-order equation (Eq. (5) in the reference Brown et al. (2015); Wingard et al. (2024)):
$\left[D_{1}^{2}+D_{2}^{2}+D_{3}^{2}+D_{4}^{2}-\frac{m^{2}c^{2}}{\hbar^{2}}\right]\psi+\frac{q}{\hbar c}(\mathbf{B}\cdot\boldsymbol{\sigma}-i\mathbf{E}\cdot\boldsymbol{\sigma})\psi=0$ (12)
where $D_{\mu}=\partial_{\mu}-\frac{iq}{\hbar c}A_{\mu}$, $\mathbf{B}=\nabla\times\mathbf{A}$, $\mathbf{E}=-\frac{1}{c}\frac{\partial\mathbf{A}}{\partial t}-\nabla\Phi$, and $\boldsymbol{\sigma}$ are the Pauli matrices.
II.2 Feshbach-Villars Linearization
The FV formalism rewrites the second-order equation as a first-order system in time by splitting the wave function :
$\psi=\phi+\chi$ (13)
$(i\frac{\partial}{\partial t}-V)\psi=m(\phi-\chi)$ (14)
This leads to two coupled equations:
$(i\frac{\partial}{\partial t}-V)(\phi+\chi)=m(\phi-\chi)$ (15)
$(i\frac{\partial}{\partial t}-V)(\phi-\chi)=\frac{p^{2}}{m}(\phi+\chi)+m(\phi+\chi)-\frac{i}{m}(\nabla V\cdot\boldsymbol{\sigma})(\phi+\chi)$ (16)
where $V=q\Phi$ is the scalar potential. These can be combined into a two-component Schrödinger-like equation:
$i\frac{\partial}{\partial t}\begin{pmatrix}\phi\\ \chi\end{pmatrix}=H\begin{pmatrix}\phi\\ \chi\end{pmatrix}$ (17)
with the FV Hamiltonian (Eq. (12) in the reference Wingard et al. (2024); Brown et al. (2015) ):
$H=\begin{pmatrix}1&1\\ -1&-1\end{pmatrix}\frac{p^{2}}{2m}+\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}m+\begin{pmatrix}1&0\\ 0&1\end{pmatrix}V-\begin{pmatrix}1&1\\ -1&-1\end{pmatrix}\frac{i}{2m}(\nabla V\cdot\boldsymbol{\sigma})$ (18)
Using Pauli matrices $\tau_{1},\tau_{2},\tau_{3}$ in the FV (component) space, the Hamiltonian is written as :
$H=(\tau_{3}+i\tau_{2})\frac{p^{2}}{2m}+\tau_{3}m+I_{2}V-(\tau_{3}+i\tau_{2})\frac{i}{2m}(\nabla V\cdot\boldsymbol{\sigma})$ (19)
[…]
II.3 FV oscillator for particles of spin-1/2
For the next discussion, to look at the DO in the FV representation, we begin with the minimal substitution $p\longrightarrow p-im\omega x$ of the momentum operator. Here $\omega$ is the oscillation frequency and $x_{\mu}=(0,x,0,0)$. Thus, by generalizing the momentum operator, the one dimensional FV Hamiltonian (19) becomes (Here $V=0$) Bouzenada et al. (2023a)
$H=(\tau_{3}+i\tau_{2})\frac{\left(p+im\omega x\right)\left(p-im\omega x\right)}{2m}+\tau_{3}m$ (20)
In this way, the Schrödinger formulation of the DO in one dimension is obtained by replacing the ansatz (13) into (15) with the following :
$\frac{1}{2m}\left[\frac{d^{2}}{dx^{2}}-m^{2}\omega^{2}x^{2}+m\omega\right]\left(\phi+\chi\right)+m\phi$ $=E\phi$ (21)
$-\frac{1}{2m}\left[\frac{d^{2}}{dx^{2}}-m^{2}\omega^{2}x^{2}+m\omega\right]\left(\phi+\chi\right)-m\chi$ $=E\chi$
After making a calculation similar to that which was made in the previous equation of the FVO, it follows that
$\left[\frac{d^{2}}{dx^{2}}-m^{2}\omega^{2}x^{2}+\left(E^{2}-m^{2}+m\omega\right)\right]\psi(x)=0$ (22)
or,
$\left[\frac{d^{2}}{dx^{2}}-\lambda_{1}x^{2}+\lambda_{2}\right]\psi(x)=0$ (23)
where we have set,
$\lambda_{1}=m^{2}\omega^{2},\quad\lambda_{2}=E^{2}-m^{2}+m\omega.$ (24)
Eq. (23) is a second order differential equation for the field $\psi$ describing the FVO dynamics in one dimensional Minkowski space-time. The solution of the above equation can be found in the literature, and a quantification condition of energy results from this solution, providing.
$\frac{E^{2}-m^{2}+m\omega}{2m}=\left(n+\frac{1}{2}\right)\omega,\quad n=0,1,2,\ldots$ (25)
Having arranged and simplified the condition (24), we find the following expression of the energy spectrum
$E=\pm\sqrt{2m\omega n+m^{2}}$ (26)
Eq. (26) shows the relativistic energy spectrum of the FVO in space-time at Minkowski, which is the same as the one obtained for the KGO in flat space-time (see Ref. Boumali (2015)). We see that this result is expected because the energy spectrum is observable and is not affected by the representation we used.
III Hamiltonian form of Dirac equation in curved spacetime
@@
The tetrad (vierbein) fields $e^{a}{}_{\mu}$ and their inverse $e_{a}{}^{\mu}$ relate the curved metric to the local Minkowski metric via $g_{\mu\nu}=e^{a}{}_{\mu}e^{b}{}_{\nu}\,\eta_{ab}$ and define the curved gamma matrices $\gamma^{\mu}=e_{a}{}^{\mu}\gamma^{a}$. We distinguish: (i) the spacetime covariant derivative $\nabla_{\mu}$ acting on tensor indices, (ii) the gauge-covariant derivative $D_{\mu}=\partial_{\mu}-iqA_{\mu}$ used for minimal electromagnetic coupling, and (iii) the spinor covariant derivative $\mathcal{D}_{\mu}=\partial_{\mu}+\frac{1}{4}\omega_{\mu}{}^{ab}\gamma_{ab}-iqA_{\mu}$, which adds the spin connection to ensure local Lorentz covariance. These derivatives coincide on scalars but differ on tensors/spinors through their respective connection terms Nakahara (2003).
@@
$\gamma_{ab}=\frac{1}{2}[\gamma_{a},\gamma_{b}],$ (30)
[…]
@@
  • $i\gamma^{0}\gamma^{i}\mathcal{D}_{i}$: Represents the kinetic energy and coupling of the spinor to the spatial curvature via the spin connection $\omega_{i}^{ab}$.
@@
  • $\frac{1}{4}\gamma^{0}\omega_{0}^{ab}\gamma_{ab}$: Represents the coupling to the spin connection in the time direction.
@@
where $\nabla_{\mu}$ is the covariant derivative compatible with the curved spacetime metric. Thus, the probability density $\rho$ is the 0-component of the current:
@@
IV Feshbach-Villars of a particle of spin-1/2 in curved spacetime: The formalism
The modified Klein-Gordon equation obtained by squaring the operator in the Dirac equation, first found by Erwin Schr"odinger Schrodinger (1932) is given by
@@ IV Feshbach-Villars of a particle of spin-1/2 in curved spacetime: The formalism
Following the procedure used by Silenko Silenko (2013) the general form of the covariant KG equation reads(39) (see Annex A)
@@
The curly bracket $\{\ldots,\ldots\}$ denotes the anticommutator. In our case, the first term is
@@
The Feshbach-Villars Metric and Inner Product
The FV transformation defined in Eq. (45) recasts the squared Dirac equation (Schr"odinger form) in Eq. (39) into a first-order Schr"odinger-type evolution equation in an enlarged state space. In particular, one introduces the two-component FV field
$\Psi_{FV}=(\phi,\chi)^{T}$
For spin–$1/2$, each component $\phi$ and $\chi$ is itself a Dirac spinor (four components in $3+1$ dimensions), so that $\Psi_{FV}$ carries eight components in total.
Unlike the standard Dirac theory, the Feshbach-Villars transformation is generally not unitary with respect to the standard positive-definite $L^{2}$ inner product. Instead, it is unitary with respect to an indefinite metric operator associated with the conserved charge of the second-order theory. In the canonical FV basis, this metric operator is defined as $\eta_{FV}=\sigma_{3}\otimes\mathbb{I}_{4}$ (and reduces to $\eta_{FV}=\sigma_{3}$ in the scalar FV-0 case):
$\eta_{FV}=\sigma_{3}\otimes\mathbb{I}_{4}=\begin{pmatrix}\mathbb{I}_{4}&0\\ 0&-\mathbb{I}_{4}\end{pmatrix}$.
The corresponding conserved inner product on a hypersurface $\Sigma_{t}$ of constant coordinate time is given by:
$\langle\Psi_{1}|\Psi_{2}\rangle_{FV}=\int_{\Sigma_{t}}d^{3}x\sqrt{h}\Psi_{1}^{\dagger}\eta_{FV}\Psi_{2},$ (47)
where $h=\det(h_{ij})$ is the determinant of the induced spatial metric. For the FV-1/2 construction based on Eq. (39), this bilinear form is the natural generalization of the scalar FV charge and is conserved by the pseudo-unitary time evolution. When the initial data are restricted to solutions of the underlying first-order Dirac equation Eq. (27), the FV charge is consistent with the usual conserved Dirac current in Eq. (35) (up to the standard identification of the positive-frequency subspace in stationary backgrounds).
Consequently, the Hamiltonian $\mathcal{H}$ derived in this framework is not Hermitian in the standard sense ($H^{\dagger}\neq H$) but rather pseudo-Hermitian with respect to $\eta_{FV}$. This condition is expressed as:
$\mathcal{H}^{\dagger}=\eta_{FV}\mathcal{H}\eta_{FV}^{-1},$
which ensures that the time evolution is pseudo-unitary, preserving the generalized inner product (the charge $Q$) rather than the standard probability norm Mostafazadeh (2002); Silenko (2008).
Equivalence with the Dirac equation and elimination of spurious solutions
Because the FV Hamiltonian is obtained from the squared Dirac operator, it is worth recalling a standard subtlety: every solution of the Dirac equation is a solution of the squared equation, but the converse is not true in general. The second-order equation admits additional (spurious) solutions that do not satisfy the original first-order constraint. In the present work, the FV formulation is derived starting from Eq. (27), and we therefore regard the FV Hamiltonian evolution as a Hamiltonian representation of Dirac dynamics on the physical subspace.
Concretely, given an FV state $\Psi_{FV}=(\phi,\chi)^{T}$, the underlying Dirac spinor is reconstructed as $\Psi=\phi+\chi$, while the auxiliary combination $\phi-\chi$ is fixed by the first-order relation in Eq. (45). If the initial data obey the Dirac constraint, then the FV evolution preserves this constraint and no spurious solutions are generated. This “Dirac restriction” is the precise sense in which the FV-1/2 formalism is equivalent to the standard covariant Dirac theory.
Using the FV transformation, the Klein–Gordon-type equation obtained by squaring the Dirac operator is rewritten as a first-order equation for $\psi$. Substituting the definitions of $\phi$ and $\chi$, we derive a coupled system of equations:
@@
$e\vec{\Sigma}\cdot\vec{B}=-e\begin{pmatrix}\vec{\sigma}&0\\ 0&\vec{\sigma}\end{pmatrix}\cdot\vec{B}.$ (70)